OpenStudy (anonymous):

Condense 1/2 log516 - 3 log5X + 4 log5Y a.log58y4/x3 b.log54 Y4/X3 c.1.5 log516Y/X d.log5(4 - x3 + Y4)

OpenStudy (campbell_st):

here are some log laws that may help 1. Addition/Multiplication \[\log _{a}m + \log _{a} n = \log _{a}(m timesn)... or... \log _{a}(mn)\] 2. Subtraction/Division \[\log _{a}m - \log _{a} n = \log _{a}(\frac{m}{n})\] 3. Power \[mlog _{a}n = \log _{a}(n^m)\] start by using the power law \[\log _{5} (16)^{\frac{1}{2}} - \log _{5}(x^3) + \log _{5}(y^4)\] \[16^\frac{1}{2} = 4\] next the addition and subtraction laws \[\log _{5}(\frac{4}{x^3})+\log _{5}y^4\] gives a final answer \[\log _{5}(\frac{4y^4}{x^3})\]